# Alleviating Distortion in Image Generation via Multi-Resolution Diffusion Models

### v1

@@ Abstract
This paper presents innovative enhancements to diffusion models by integrating a novel multi-resolution network and time-dependent layer normalization. Diffusion models have gained prominence for their effectiveness in high-fidelity image generation. While conventional approaches rely on convolutional U-Net architectures, recent Transformer-based designs have demonstrated superior performance and scalability. However, Transformer architectures, which tokenize input data (via “patchification”), face a trade-off between visual fidelity and computational complexity due to the quadratic nature of self-attention operations concerning token length. While larger patch sizes enable attention computation efficiency, they struggle to capture fine-grained visual details, leading to image distortions. To address this challenge, we propose augmenting the Diffusion model with the Multi-Resolution network (DiMR), a framework that refines features across multiple resolutions, progressively enhancing detail from low to high resolution. Additionally, we introduce Time-Dependent Layer Normalization (TD-LN), a parameter-efficient approach that incorporates time-dependent parameters into layer normalization to inject time information and achieve superior performance. Our method’s efficacy is demonstrated on the class-conditional ImageNet generation benchmark, where DiMR-XL variants outperform prior diffusion models, setting new state-of-the-art FID scores of 1.70 on ImageNet $256\times 256$ and 2.89 on ImageNet $512\times 512$.
@@ 1 Introduction
To demonstrate its effectiveness, we evaluate DiMR on the class-conditional ImageNet generation benchmark [7]. On ImageNet $64\times 64$, DiMR-M (133M parameters) and DiMR-L (284M), without classifier-free guidance [18], achieve FID scores of 3.65 and 2.21, respectively, outperforming the Transformer-based U-ViT-M/4 and U-ViT-L/4 by 2.20 and 2.05 FID. On ImageNet $256\times 256$, DiMR-XL (505M) achieves FID scores of 4.50 and 1.70, without and with classifier-free guidance, respectively. On ImageNet $512\times 512$, DiMR-XL (525M) achieves FID scores of 7.93 and 2.89, without and with classifier-free guidance, respectively. These results demonstrate superior performance compared to all previous methods, despite having similar or smaller model sizes, establishing a new state-of-the-art performance. In summary, our main contributions are as follows:
@@ 2 Related Work
Architecture for diffusion models. Early diffusion models employed convolutional U-Net architectures [46] as the denoising network, which were subsequently strengthened through explorations of either computing attention [8, 38] or performing diffusion directly at multiple scales [20, 13]. Recently, Transformer-based architectures [3, 40, 14] along with other explorations [62, 56] have emerged as promising alternatives, showcasing superior performance and scalability. Specifically, for Transformer-based architectures, U-ViT [3] treats all inputs, including time, condition, and noisy image patches, as tokens and employs long-skip connections between shallow and deep transformer layers inspired by U-Net. Similarly, DiT [40] leverages Vision Transformers (ViTs) [9] to systematically explore the design space under the Latent Diffusion Models (LDMs) framework, demonstrating favorable properties such as scalability, robustness, and efficiency. In this study, we introduce the Multi-Resolution Network as a new denoising architecture for diffusion models, featuring a multi-branch design where each branch is dedicated to processing a specific resolution.
@@ 4.2 Time-Dependent Layer Normalization
Unlike adaLN, which learns additional re-scaling $\gamma_{1}$ and re-centering $\beta_{1}$ variables, TD-LN directly incorporates the time-dependent $\gamma(t)$ and $\beta(t)$ into layer normalization, eliminating the need for a parameter-heavy MLP. Furthermore, TD-LN is a versatile mechanism, enabling the injection of time information into both Transformer blocks and ConvNeXt blocks. In DiMR, we replace all layer normalizations with the proposed TD-LN, and treat the class condition $c$ as input token for the Transformer blocks (ConvNeXt blocks do not take class condition token).
@@ 4.4 DiMR Model Variants
We now introduce the DiMR model variants, scaled appropriately for different model sizes. We present three sizes: DiMR-M (medium, 133M parameters), DiMR-L (large, 284M parameters), and DiMR-XL (extra-large, around 500M parameters). Three hyperparameters—$R$ (number of branches), $N$ (number of layers per branch), and $D$ (hidden size per branch)—define each DiMR variant. Specifically, $R$ determines the number of branches in the multi-resolution network. We append 2R or 3R to the model name to indicate whether two or three branches are used. The number of layers $N$ in the multi-resolution network is represented as a tuple of $R$ numbers, where the $r$-th number specifies the number of layers in the $r$-th branch. Similarly, the hidden size $D$ is also a tuple of $R$ numbers. The model variants details are presented in Tab. 3 in Sec. B in the Appendix.
@@ 5.1 Experimental Setup
Implementation details. For a fair comparison with DiT [40], we train the $256\times 256$ and $512\times 512$ models for 1M iterations, and also report results for 500K iterations to compare with U-ViT [3]. We train the $64\times 64$ models for 300K iterations, following the U-ViT protocol. Our training hyperparameters are almost entirely retained from U-ViT [3]. We did not tune learning rates, decay/warm-up schedules, Adam $\beta_{1}$/$\beta_{2}$ values, or weight decays. Further details on hyperparameters and configurations are provided in Sec. C in the Appendix.
@@ 5.2 State-of-the-Art Diffusion Models
Table 1: Class-conditional image generation on ImageNet $256\times 256$ and ImageNet $512\times 512$. We report training epochs, number of parameters (#Params), GFLOPs, and FID-50K with and without Classifier-Free Guidance (CFG). Best results are marked in bold. Our DiMR achieves state-of-the-art results, and even surpasses most recent large-scale models on ImageNet $256\times 256$ (marked in gray).
@@
ImageNet $256\times 256$. From Tab. 1a, we observe that our DiMR outperforms all previous diffusion-based models and achieves a state-of-the-art FID-50K score of 1.70. Specifically, with a comparable model size and equal or fewer training epochs, our model surpasses previous state-of-the-art transformer-based diffusion models, including U-ViT[3] (1.77 vs. 2.29 with Classifier-Free Guidance [18] (CFG) and 4.87 vs. 6.58 without CFG) and DiT [40] (1.70 vs. 2.27 with CFG and 4.50 vs. 9.62 without CFG). Moreover, with only 505M parameters, our model even outperforms the most recent large-scale diffusion model (Large-DiT [1], 3.0B) by a significant margin (1.77 vs. 2.10), indicating the strong capabilities of the proposed DiMR.
ImageNet $512\times 512$. Our DiMR outperforms all previous diffusion-based models on ImageNet $512\times 512$ and achieves a state-of-the-art FID-50K score of 2.89 as shown in Tab. 1b. It is worth noting that, although both Gflops and model sizes are critical for improving performance, as discussed in the DiT paper [40], we still outperform it with only $39.2\%$ of the GFLOPs and $77.8\%$ of the model size, improving the FID-50K from 3.04 to 2.89.
@@ 5.3 Alleviating Distortion
5.4 Ablation Studies
We conduct ablation experiments on ImageNet $64\times 64$, step by step starting from the baseline U-ViT-M/4 [3], to verify the effectiveness of the proposed designs, as presented in Tab. 2.
@@ Appendix B DiMR Model Variants
We introduce the DiMR model variants, scaled appropriately for different model sizes. We present three sizes: DiMR-M (medium, 132.7M parameters), DiMR-L (large, 284.0M parameters), and DiMR-XL (extra-large, around 500M parameters). Three hyperparameters—$R$ (number of branches), $N$ (number of layers per branch), and $D$ (hidden size per branch)—define each DiMR variant. Specifically, $R$ determines the number of branches in the multi-resolution network. We append 2R or 3R to the model name to indicate whether two or three branches are used. The number of layers $N$ in the multi-resolution network is represented as a tuple of $R$ numbers, where the $r$-th number specifies the number of layers in the $r$-th branch. Similarly, the hidden size $D$ is also a tuple of $R$ numbers. We follow a straightforward scaling rule: most layers are stacked in the first branch, which is processed by Transformer blocks, while the remaining branches use only half the number of layers of the first branch. Additionally, when the resolution is doubled, the hidden size is reduced by a factor of two. The model variants are illustrated in Tab. 3.
@@ Appendix C Implementation Details
Model DiMR-M/3R DiMR-L/3R DiMR-XL/2R DiMR-XL/3R
Parameters 133M 284M 505M 525M
Image Resolution $64\times 64$ $64\times 64$ $256\times 256$ $512\times 512$
@@
Latent shape - - $32\times 32\times 4$ $64\times 64\times 4$
Image decoder - - sd-vae-ft-ema sd-vae-ft-ema
Number of branches $R$ 3 3 2 3
Blocks in 1st branch Transformer Transformer Transformer Transformer
    # Layers 15 33 39 39
    # Dimensions 768 768 960 960
    # Heads 12 12 16 16
    Resolution $16\times 16$ $16\times 16$ $16\times 16$ $16\times 16$
    Loss Coeffs. $1/16$ $1/16$ $1/4$ $1/16$
Blocks in 2nd branch ConvNeXt ConvNeXt ConvNeXt ConvNeXt
    # Layers 8 17 20 20
    # Dimensions 384 384 480 480
    Kernel size $7\times 7$ $7\times 7$ $7\times 7$ $7\times 7$
    Resolution $32\times 32$ $32\times 32$ $32\times 32$ $32\times 32$
    Loss Coeffs. $1/4$ $1/4$ $1$ $1/4$
Blocks in 3rd branch ConvNeXt ConvNeXt - ConvNeXt
    # Layers 8 17 - 20
    # Dimensions 192 192 - 240
@@
Batch size 1024 1024 1024 1024
Training iterations 300K 300K 1M 1M
Warm-up steps 5K 5K 5K 5K
Optimizer AdamW AdamW AdamW AdamW
learning rate $3\times 10^{-4}$ $3\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Weight decay $0.03$ $0.03$ $0.03$ $0.03$
Betas $(0.99,0.99)$ $(0.99,0.99)$ $(0.99,0.99)$ $(0.99,0.99)$
Training devices 8 A100 16 A100 16 A100 16 A100
Training time 62 hours 80 hours 172 hours 288 hours
Sampler DPM-Solver DPM-Solver DPM-Solver DPM-Solver
Sampling steps 50 50 250 250
@@ Appendix D Additional Experimental Results
MDT-G [11] Mask. + Diff. 1299 700M 121 1.79 283.0 0.81 0.61
@@
ImageNet $256\times 256$. We compare DiMR with state-of-the-art generative models on ImageNet $256\times 256$ with classifier-free guidance in Tab. 6. Compared to U-ViT [3] in a fair setting, our DiMR-XL/2R with 505M parameters, trained for 400 epochs, significantly outperforms U-ViT-H/2 with 501M parameters, also trained for 400 epochs, by 0.52 in FID and 21.8 in IS. In comparison with the recently popular diffusion model DiT [40], DiMR-XL/2R trained for 800 epochs consistently outperforms DiT-L/2 with 458M parameters by 3.32 in FID and 121.8 in IS. DiMR-XL/2R even surpasses the larger variant DiT-XL/2 with 675M parameters by 0.57 in FID and 11.8 in IS. Notably, DiT models require training for 1399 epochs, while DiMR-XL/2R achieves superior performance with only 800 epochs. As a result, DiMR-XL/2R sets a new state-of-the-art on ImageNet $256\times 256$ image generation with an FID of 1.70 and an IS of 289.0.
ImageNet $512\times 512$. We compare DiMR with state-of-the-art generative models on ImageNet $512\times 512$ with classifier-free guidance in Tab. 7. Under the same training setting for 400 epochs, our DiMR-XL/3R with 525M parameters outperforms U-ViT-H/4 with 501M parameters by 0.82 in FID and 21.3 in IS. Compared to DiT-XL/2 with 675M parameters, DiMR-XL/3R shows slight improvements of 0.15 in FID and 49.0 in IS. Consequently, DiMR-XL/3R sets the state-of-the-art performance on ImageNet $512\times 512$ image generation when using diffusion-based generative models.

### v2

@@ Abstract
This paper presents innovative enhancements to diffusion models by integrating a novel multi-resolution network and time-dependent layer normalization. Diffusion models have gained prominence for their effectiveness in high-fidelity image generation. While conventional approaches rely on convolutional U-Net architectures, recent Transformer-based designs have demonstrated superior performance and scalability. However, Transformer architectures, which tokenize input data (via “patchification”), face a trade-off between visual fidelity and computational complexity due to the quadratic nature of self-attention operations concerning token length. While larger patch sizes enable attention computation efficiency, they struggle to capture fine-grained visual details, leading to image distortions. To address this challenge, we propose augmenting the Diffusion model with the Multi-Resolution network (DiMR), a framework that refines features across multiple resolutions, progressively enhancing detail from low to high resolution. Additionally, we introduce Time-Dependent Layer Normalization (TD-LN), a parameter-efficient approach that incorporates time-dependent parameters into layer normalization to inject time information and achieve superior performance. Our method’s efficacy is demonstrated on the class-conditional ImageNet generation benchmark, where DiMR-XL variants surpass previous diffusion models, achieving FID scores of 1.70 on ImageNet $256\times 256$ and 2.89 on ImageNet $512\times 512$. Our best variant, DiMR-G, further establishes a state-of-the-art 1.63 FID on ImageNet $256\times 256$.
@@ 1 Introduction
To demonstrate its effectiveness, we evaluate DiMR on the class-conditional ImageNet generation benchmark [6]. On ImageNet $64\times 64$, DiMR-M (133M parameters) and DiMR-L (284M), without classifier-free guidance [18], achieve FID scores of 3.65 and 2.21, respectively, outperforming the Transformer-based U-ViT-M/4 and U-ViT-L/4 by 2.20 and 2.05 FID. On ImageNet $256\times 256$, DiMR-XL (505M) achieves FID scores of 4.50 without classifier-free guidance and 1.70 with classifier-free guidance. Meanwhile, DiMR-G (1.06B) further improves the FID scores to 3.56 without classifier-free guidance and 1.63 with classifier-free guidance. On ImageNet $512\times 512$, DiMR-XL (525M) achieves FID scores of 7.93 and 2.89, without and with classifier-free guidance, respectively. These results demonstrate superior performance compared to all previous methods, despite having similar or smaller model sizes, establishing a new state-of-the-art performance. In summary, our main contributions are as follows:
@@ 2 Related Work
Architecture for diffusion models. Early diffusion models employed convolutional U-Net architectures [47] as the denoising network, which were subsequently strengthened through explorations of either computing attention [7, 39] or performing diffusion directly at multiple scales [20, 13]. Recently, Transformer-based architectures [2, 41, 14] along with other explorations [64, 57, 27] have emerged as promising alternatives, showcasing superior performance and scalability. Specifically, for Transformer-based architectures, U-ViT [2] treats all inputs, including time, condition, and noisy image patches, as tokens and employs long-skip connections between shallow and deep transformer layers inspired by U-Net. Similarly, DiT [41] leverages Vision Transformers (ViTs) [8] to systematically explore the design space under the Latent Diffusion Models (LDMs) framework, demonstrating favorable properties such as scalability, robustness, and efficiency. In this study, we introduce the Multi-Resolution Network as a new denoising architecture for diffusion models, featuring a multi-branch design where each branch is dedicated to processing a specific resolution.
@@ 3 Preliminary
Classifier-free guidance [18], an effective approach to generating high-fidelity samples, combines the score estimates from a conditional diffusion model and a jointly trained unconditional diffusion model. Formally, the classifier-free guidance encourages the sampled $\bm{x}$ to have high $p(\bm{x}|c)$ by setting: $\hat{\epsilon}_{\theta}(\bm{x}_{t},c)=\epsilon_{\theta}(\bm{x}_{t},\emptyset)+s\cdot\nabla_{\bm{x}}\log p(\bm{x}|c)\propto\epsilon_{\theta}(\bm{x}_{t},\emptyset)+s\cdot(\epsilon_{\theta}(\bm{x}_{t},c)-\epsilon_{\theta}(\bm{x}_{t},\emptyset))$, where $s$ is the scale of guidance, $s\geq 1$, and setting $s=1$ becomes the standard sampling. Following prior arts [2, 41], we also exploit classifier-free guidance.
@@ 4.2 Time-Dependent Layer Normalization
Unlike adaLN, which learns additional re-scaling $\gamma_{1}$ and re-centering $\beta_{1}$ variables, TD-LN directly incorporates the time-dependent $\gamma(t)$ and $\beta(t)$ into layer normalization, eliminating the need for a parameter-heavy MLP. Furthermore, TD-LN is a versatile mechanism, enabling the injection of time information into both Transformer blocks and ConvNeXt blocks. In DiMR, we replace all layer normalizations with the proposed TD-LN, and treat the class condition $c$ as input token for the Transformer blocks.
@@ 4.4 DiMR Model Variants
We now introduce the DiMR model variants, scaled appropriately for different model sizes. We present four sizes: DiMR-M (medium, 133M parameters), DiMR-L (large, 284M parameters), DiMR-XL (extra-large, around 500M parameters) and DiMR-G (giant, 1.06B parameters). Three hyperparameters—$R$ (number of branches), $N$ (number of layers per branch), and $D$ (hidden size per branch)—define each DiMR variant. Specifically, $R$ determines the number of branches in the multi-resolution network. We append 2R or 3R to the model name to indicate whether two or three branches are used. The number of layers $N$ in the multi-resolution network is represented as a tuple of $R$ numbers, where the $r$-th number specifies the number of layers in the $r$-th branch. Similarly, the hidden size $D$ is also a tuple of $R$ numbers. We follow a straightforward scaling rule: most layers are stacked in the first branch, which is processed by Transformer blocks, while the remaining branches use only half the number of layers of the first branch. Additionally, when the resolution is doubled, the hidden size is reduced by a factor of two. The model variants details are presented in Tab. 4 in Sec. B in the Appendix.
@@ 5.1 Experimental Setup
Implementation details. We use AdamW optimizer [36] with a constant learning rate of $2\times 10^{-4}$ for most experiments, except for the $64\times 64$ models where we use $3\times 10^{-4}$. A batch size of 1024 is used for most architectures. For a fair comparison with DiT [41], we train the $256\times 256$ and $512\times 512$ models for 1M iterations, and also report results for 500K iterations to compare with U-ViT [2]. We train the $64\times 64$ models for 300K iterations, following the U-ViT protocol. Our training hyperparameters are almost entirely retained from U-ViT [2]. We did not tune learning rates, decay/warm-up schedules, Adam $\beta_{1}$/$\beta_{2}$ values, or weight decays. Further details on hyperparameters and configurations are provided in Sec. C in the Appendix.
@@ 5.2 State-of-the-Art Diffusion Models
Table 1: Class-conditional image generation on ImageNet $256\times 256$ and ImageNet $512\times 512$. We report training epochs, number of parameters (#Params), GFLOPs, and FID-50K with and without Classifier-Free Guidance (CFG). Best results are marked in bold.
@@
ImageNet $256\times 256$. From Tab. 1a, we observe that our DiMR-XL/2R outperforms all previous diffusion-based models and achieves a state-of-the-art FID-50K score of 1.70. Specifically, with a comparable model size and equal or fewer training epochs, our model surpasses previous state-of-the-art transformer-based diffusion models, including U-ViT[2] (1.77 vs. 2.29 with Classifier-Free Guidance [18] (CFG) and 4.87 vs. 6.58 without CFG) and DiT [41] (1.70 vs. 2.27 with CFG and 4.50 vs. 9.62 without CFG). Our best model, DiMR-G/2R, scales up to the billion-parameter level, setting a new state-of-the-art with an FID of 1.63 with CFG and 3.56 without CFG.
ImageNet $512\times 512$. Our DiMR outperforms all previous diffusion-based models on ImageNet $512\times 512$ and achieves a state-of-the-art FID-50K score of 2.89 as shown in Tab. 1b. It is worth noting that, although both Gflops and model sizes are critical for improving performance, as discussed in the DiT paper [41], we still outperform it with only $39.2\%$ of the GFLOPs and $77.8\%$ of the model size, improving the FID-50K from 3.04 to 2.89. As transformers and diffusion models have demonstrated good scaling behavior, we believe that further scaling up our DiMR will lead to better performance, which we have left as future work.
@@ 5.3 Alleviating Distortion
5.4 Ablation Studies
We conduct the primary ablation experiments on ImageNet $64\times 64$, progressively building on the baseline U-ViT-M/4 [2] to validate the effectiveness of the proposed designs, leading to our final model, DiMR-M/3R, as presented in Tab. 2. Additionally, we explore alternative design choices on ImageNet $256\times 256$ with DiMR-XL/2R, including adopting a pure convolutional architecture, replacing addition with concatenation in feature cascading, and introducing skip connections between branches, as shown in Tab. 3.
Transformer is essential for low-resolution processing. As shown in Table 3a, replacing the Transformer blocks in the 1st (lowest-resolution) branch with ConvNeXt blocks results in a DiMR variant that uses only convolutional layers. However, this configuration performs worse compared to combining Transformer blocks with ConvNeXt blocks across different resolutions. This indicates that Transformers are more effective at capturing fine-grained details, while their usage at the lowest resolution maintains a manageable computational cost.
Simple addition suffices for multi-resolution feature cascading. As shown in Table 3b, a straightforward addition operation effectively transfers information from lower-resolution features to higher-resolution features. Replacing addition with concatenation leads to slightly worse results. We also validate the necessity of adding skip-connection between branches. As shown in Table 3c, introducing skip-connection not only degrades performance but also complicates the model architecture. Therefore, we adopt a simple upsampling followed by an addition operation for feature cascading.
Table 3: Design choices. We empirically experiment with different design choices in model architecture (Tab 3a), feature cascade (Tab 3b) and skip-connection (Tab 3c). We report FID-50K scores with classifier-free guidance (CFG) after 400 training epochs.
[…]
DiMR-XL/2R ConvNeXt 2.09
Transformer 1.77
(a) Pure Conv v.s. Hybrid
[…]
DiMR-XL/2R Concatenation 2.06
Addition 1.77
(b) Concatenation v.s. Addition
[…]
DiMR-XL/2R 1.96
1.77
(c) w/ v.s. w/o Skip-Connection
@@ Appendix B DiMR Model Variants
We introduce the DiMR model variants, scaled appropriately for different model sizes. We present four sizes: DiMR-M (medium, 132.7M parameters), DiMR-L (large, 284.0M parameters), DiMR-XL (extra-large, around 500M parameters), and DiMR-G (giant, 1.06B parameters) Three hyperparameters—$R$ (number of branches), $N$ (number of layers per branch), and $D$ (hidden size per branch)—define each DiMR variant. Specifically, $R$ determines the number of branches in the multi-resolution network. We append 2R or 3R to the model name to indicate whether two or three branches are used. The number of layers $N$ in the multi-resolution network is represented as a tuple of $R$ numbers, where the $r$-th number specifies the number of layers in the $r$-th branch. Similarly, the hidden size $D$ is also a tuple of $R$ numbers. We follow a straightforward scaling rule: most layers are stacked in the first branch, which is processed by Transformer blocks, while the remaining branches use only half the number of layers of the first branch. Additionally, when the resolution is doubled, the hidden size is reduced by a factor of two. The model variants are illustrated in Tab. 4.
@@ Appendix C Implementation Details
Model DiMR-M/3R DiMR-L/3R DiMR-XL/2R DiMR-XL/3R DiMR-G/2R
Parameters 133M 284M 505M 525M 1.06B
Image Resolution $64\times 64$ $64\times 64$ $256\times 256$ $512\times 512$ $256\times 256$
@@
Latent shape - - $32\times 32\times 4$ $64\times 64\times 4$ $32\times 32\times 4$
Image decoder - - sd-vae-ft-ema sd-vae-ft-ema sd-vae-ft-ema
Number of branches $R$ 3 3 2 3 2
Blocks in 1st branch Transformer Transformer Transformer Transformer Transformer
    # Layers 15 33 39 39 57
    # Dimensions 768 768 960 960 1152
    # Heads 12 12 16 16 16
    Resolution $16\times 16$ $16\times 16$ $16\times 16$ $16\times 16$ $16\times 16$
    Loss Coeffs. $1/16$ $1/16$ $1/4$ $1/16$ $1/4$
Blocks in 2nd branch ConvNeXt ConvNeXt ConvNeXt ConvNeXt ConvNeXt
    # Layers 8 17 20 20 29
    # Dimensions 384 384 480 480 576
    Kernel size $7\times 7$ $7\times 7$ $7\times 7$ $7\times 7$ $7\times 7$
    Resolution $32\times 32$ $32\times 32$ $32\times 32$ $32\times 32$ $32\times 32$
    Loss Coeffs. $1/4$ $1/4$ $1$ $1/4$ $1$
Blocks in 3rd branch ConvNeXt ConvNeXt - ConvNeXt -
    # Layers 8 17 - 20 -
    # Dimensions 192 192 - 240 -
@@
Batch size 1024 1024 1024 1024 1024
Training iterations 300K 300K 1M 1M 1M
Warm-up steps 5K 5K 5K 5K 5K
Optimizer AdamW AdamW AdamW AdamW AdamW
learning rate $3\times 10^{-4}$ $3\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Weight decay $0.03$ $0.03$ $0.03$ $0.03$ $0.03$
Betas $(0.99,0.99)$ $(0.99,0.99)$ $(0.99,0.99)$ $(0.99,0.99)$ $(0.99,0.99)$
Training devices 8 A100 16 A100 16 A100 16 A100 32 A100
Training time 62 hours 80 hours 172 hours 288 hours 400 hours
Sampler DPM-Solver DPM-Solver DPM-Solver DPM-Solver DPM-Solver
Sampling steps 50 50 250 250 250
@@ Appendix D Additional Experimental Results
TiTok-S-128 [66] Mask. 800 287M - 1.97 281.8 - -
MDT-G [10] Mask. + Diff. 1299 676M 119 1.79 283.0 0.81 0.61
MDTv2-G [11] Mask. + Diff. 919 675M 119 1.58 314.7 0.79 0.65
MaskBit [60] Mask. 1080 305M - 1.52 328.6 - -
@@
ImageNet $256\times 256$. We compare DiMR with state-of-the-art generative models on ImageNet $256\times 256$ with classifier-free guidance in Tab. 7. Compared to U-ViT [2] in a fair setting, our DiMR-XL/2R with 505M parameters, trained for 400 epochs, significantly outperforms U-ViT-H/2 with 501M parameters, also trained for 400 epochs, by 0.52 in FID and 21.8 in IS. In comparison with the recently popular diffusion model DiT [41], DiMR-XL/2R trained for 800 epochs consistently outperforms DiT-L/2 with 458M parameters by 3.32 in FID and 121.8 in IS. DiMR-XL/2R even surpasses the larger variant DiT-XL/2 with 675M parameters by 0.57 in FID and 11.8 in IS. Notably, DiT models require training for 1399 epochs, while DiMR-XL/2R achieves superior performance with only 800 epochs. Furthermore, scaling up to DiMR-G/2R sets a new state-of-the-art for ImageNet $256\times 256$ image generation, achieving an FID of 1.63 and an IS of 292.5.
ImageNet $512\times 512$. We compare DiMR with state-of-the-art generative models on ImageNet $512\times 512$ with classifier-free guidance in Tab. 8. Under the same training setting for 400 epochs, DiMR-XL/3R with 525M parameters outperforms U-ViT-H/4 with 501M parameters by 0.82 in FID and 21.3 in IS. Compared to DiT-XL/2 with 675M parameters, DiMR-XL/3R shows slight improvements of 0.15 in FID and 49.0 in IS. Overall, DiMR-XL/3R demonstrates performance comparable to other state-of-the-art generative models for ImageNet $512\times 512$ image generation.